OpenStudy (anonymous):

please help me to solve this inequality 6 is greater than or equal to - 4/5n

OpenStudy (anonymous):

\[6 \geq -\frac{ 4 }{ 5n } \]? or . . .\[6 \geq -\frac{ 4 }{ 5 }n\] ?

OpenStudy (anonymous):

\[6\ge-\frac{ 4 }{ 5 }n\]

OpenStudy (anonymous):

first multiply both sides by -5/4 that will get 'n' by itself. and because you are multiplying by a negative, don't forget to switch the sign

OpenStudy (anonymous):

so \[n \le-\frac{ 24 }{ 30 }\]

OpenStudy (anonymous):

when you multiply fractions, you multiply the numerators across, and also the denominators .. like this: \[6 * -\frac{ 5 }{ 4 } = \frac{ 6 }{ 1 } * -\frac{ 5 }{ 4 } = \frac{ 6 * (-5) }{ 1 * 4 } = \frac{ -30 }{ 4 } = . . .\]

OpenStudy (anonymous):

oh give me a minute

OpenStudy (anonymous):

so \[n \le-\frac{ 30 }{ 4 }\] sorry im confused

OpenStudy (anonymous):